A watch a second hand 2 cm long. Calculate :i) the speed of the tip of the second handsii) the velocity of the tip at t=0 and t=15 seconds.iii) change in velocity from t=0 to t=15 seconds.

iv)- Average acceleration


1
Expert's answer
2021-08-26T08:10:44-0400

Given:

"r=0.02\:\rm m"

"T=60\:\rm s"


(i) the speed of the tip of the second hand

"v=\frac{2\pi r}{T}=\frac{2\pi*0.02}{60}=0.0021\:\rm m/s=2.1\: mm/s"

ii) the velocity of the tip at  "t=0"


"v_1=2.1\:\rm mm/s\quad to\: the\;right"

and "t=15" seconds

"v_2=2.1\:\rm mm/s\quad downward"

iii) change in velocity from "t=0"  to "t=15"  seconds

"\Delta v=\sqrt{2}v=\sqrt{2}*2.1\:\rm mm/s=3.0\:\rm mm/s"

iv) the average acceleration

"a=\frac{\Delta v}{\Delta t}=\frac{3.0\:\rm mm/s}{15\:\rm s}=0.2\:\rm mm/s^2"